# RoboTransfer: Geometry-Consistent Video Diffusion for Robotic Visual Policy Transfer

### v1

@@ 4.1 Synthesis Quality
4.2 Real Robot Experiments
@@
To evaluate RoboTransfer, we conducted experiments on a real-world robotic platform, assessing how synthetic data impacts visual policy generalization across varied environments.
@@
Experimental Platform and Setup Details. Our experiments were conducted on the Agilex Cobot Magic platform, equipped with two PIPER robotic arms and three Intel RealSense D435i cameras—two mounted on the wrists and one positioned overhead. We used only RGB data for policy learning, creating a challenging visual understanding problem while reflecting realistic deployment constraints.
Benchmark Task and Evaluation Methodology. We selected a dual-arm manipulation sequence with four phases: (i) left arm grasps a spoon; (ii) spoon is placed on a tray; (iii) right arm grasps the spoon; (iv) spoon is placed in a cup. We evaluated across three test conditions: identical objects/background (Same), novel object appearances (Diff-Obj), and variations in both objects and environment (Diff-All). Testing protocols are detailed in the supplementary material. We introduced a Stage Score (0-4) corresponding to successfully completed phases, providing finer-grained insights than binary success/failure metrics. This reveals how far policies progress through the task sequence, particularly valuable when analyzing generalization.
Implementation Details. We adopted ACT [67] as our baseline architecture without modifications. For each task, we collected 100 expert demonstrations via ALOHA [68] teleoperation. Our methodology follows a pretrain+finetune paradigm. We trained baseline policies on real data for 100k steps, then fine-tuned with synthetic data for 50k additional steps (learning rate reduced to 1/10). For the grasp spoon task, training on real data required 8 NVIDIA H20 GPUs for approximately 24 hours, while the synthetic data fine-tuning phase utilized the same hardware configuration for about 12 hours. Additional details of the training protocol are given in the supplementary material.
Effectiveness of Synthetic Data Augmentation.
Table 3: Performance comparison across different test conditions and models. Success Rate (%) and Stage Score (0-4) are reported for each configuration. The best results for each test conditions are highlighted in bold.
Model Same Diff-Obj Diff-All
Success (%) Score Success (%) Score Success (%) Score
Baseline (Real only) 80.0 3.6 33.3 2.7 13.3 1.6
+ Obj Augmentation 44.4 3.0 22.2 2.0
+ Obj+Bg Augmentation 46.7 3.0
Table 3 presents the comparative results. Results show significant performance degradation when baseline policies face novel visual contexts, with success rates dropping from 80.0% to 13.3%. However, policies fine-tuned with our generated data show considerable resilience. In Diff-Obj conditions, success rates improved by $33.3\%$ (relative), while in Diff-All scenarios, we observed a $251\%$ relative improvement. Stage Score results further confirm that augmented models make more progress through task phases, indicating more robust manipulation understanding even when tasks aren’t fully completed.
Effect of Synthetic Data Proportions. We investigated how different proportions of synthetic data affect performance by evaluating synthetic ratios of 25%, 50%, 75%, and 100%. Figure 7 shows these results for the Diff-All conditions. Performance follows a non-linear relationship with synthetic data proportion, peaking at 50% and showing diminishing returns beyond. Success rates increased from 13.3% with no synthetic data to 46.7% at 50% synthetic ratio—a 3.5x improvement. Stage scores similarly peaked at 3.0 with 50% mixture versus 1.6 without synthetic data. Notably, using exclusively synthetic data (100%) yielded 40.0% success rate and 2.9 stage score—improvements over purely real data but below the balanced 50% mixture. This confirms complementary contributions: synthetic data provides diversity while real data contributes physical fidelity, with optimal interaction at equal proportions. RoboTransfer preserves geometric consistency while enabling systematic appearance variation, addressing a fundamental challenge in robot learning. Additional details on model configuration and extended evaluations are in the supplementary material.
@@
Figure 7: Performance metrics across synthetic data mixing ratios. Left: Success rate vs. ratio. Right: Average stage score vs. ratio. Note that 0% represents our baseline model trained only on real data. Both metrics peak at 50% synthetic data, suggesting an optimal balance between synthetic diversity and real-world physical fidelity.
@@ 5 Conclusion
In this work, we proposed RoboTransfer, a novel diffusion-based data synthesis framework in robotics, which integrates multi-view geometry while offering explicit control over background and object attributes. We also introduced our dataset processing pipeline, which provides triplets incorporating both global geometry and appearance control conditions. The experiments and evaluations show that our RoboTransfer can produce multi-view consistency data for training robotic manipulation policies. Therefore, the generalization of visuomotor policy gets significantly improved.
Limitation. Although the method is general and flexible, several limitations remain that we aim to address in future work. First, our approach currently focuses on augmenting the appearance of objects and scenes, rather than generating new motions. Moving forward, we plan to integrate RoboTransfer with additional simulators to broaden its applicability. Furthermore, we aim to scale up both the size and diversity of the training data, which will enhance the variety and robustness of the generated outputs.
Appendix
@@
1
Appendix A Dataset Construction Details
For the training of RoboTransfer, we collected data and designed a dedicated dataset construction pipeline (see Figure 3). In this section, we provide additional implementation details.
A.1 Data Collection
Robot Platform: We construct a large-scale robotic demonstration dataset using the Agilex Cobot Magic platform, adhering to standardized data collection protocols established in prior work [44, 39]. The robot is equipped with two PIPER robotic arms. As illustrated in Figure 8, each demonstration includes synchronized RGB-D streams captured from three Intel RealSense D435i cameras: one mounted at two hand-eye viewpoint and one positioned overhead.
Dataset Design: To capture a broad spectrum of manipulation scenarios, we systematically vary tasks, backgrounds, and manipulable objects. Specifically, we design twelve distinct manipulation tasks (visualized in Figure 9), each representing diverse interaction patterns. For each task, we collect 100 demonstration segments across 10 unique object configurations, resulting in a total of 1,000 samples per task. The dataset incorporates a variety of backgrounds, including textured tabletops and cluttered surfaces, to reflect realistic operating conditions. Manipulable objects are diversified in terms of shape, size, and material, enhancing both the visual richness and physical complexity of the dataset. This comprehensive design ensures a high degree of diversity and realism, which is crucial for training robust and generalizable visual policies.
@@ A.1 Data Collection
Figure 8: Visualizations of robot platform.
@@
Figure 9: Visualizations of the dataset.
A.2 Geometry conditions
Depth conditions: To enforce both spatial and temporal consistency, we incorporate geometry-conditioned supervision into our dataset construction pipeline. During data collection, we obtain synchronized RGB-D streams, including multi-view depth maps. However, depth maps from the RGB-D sensors are often noisy and incomplete, particularly in regions with low reflectivity (e.g., black surfaces), leading to domain gaps across samples. In contrast, depth maps rendered from simulators are typically complete and noise-free.
While we explored existing depth completion methods, we found that many suffer from limited temporal consistency and poor generalization across diverse scenes. To address these limitations, we adopt the state-of-the-art Video Depth Anything (VDA) model [9], which produces temporally coherent and spatially complete depth maps.
Despite the high-quality temporal consistency of depth estimates from Video Depth Anything (VDA), these predictions lack a globally consistent scale across multiple views. To address this, we align the VDA-predicted depth maps with the RGB-D sensor measurements via a robust least-squares scale fitting procedure. While standard least-squares regression offers a closed-form solution for affine scale alignment, it is susceptible to errors introduced by sensor noise and depth outliers. To mitigate this, we introduce a robust multi-frame fitting strategy (Figure 10) that aggregates depth correspondences across time and iteratively filters outliers based on residual error thresholds. This iterative refinement ensures stable and accurate recovery of the global scale transformation. Our approach effectively preserves the geometric completeness provided by VDA while anchoring depth predictions to the real-world scale captured by the RGB-D sensor. The resulting alignment ensures both spatial accuracy across views and temporal smoothness across frames, enabling reliable depth conditions.
Input:
Predicted Depths $\mathbf{D}_{\text{pred}}\in\mathbb{R}^{B\times H\times W}$,
Sensor Depths $\mathbf{D}_{\text{sensor}}\in\mathbb{R}^{B\times H\times W}$,
Output:
Scaled Depths $\mathbf{D}_{\text{scaled}}\in\mathbb{R}^{B\times H\times W}$,
1
2Initialize Mask: $\mathcal{M}\leftarrow(\mathbf{D}_{\text{sensor}}>\epsilon)\land(\mathbf{D}_{\text{pred}}>\epsilon)$ ;
3
4for 2 iterations do
5       $\mathcal{E}\leftarrow|\mathbf{D}_{\text{pred}}-\mathbf{D}_{\text{sensor}}|\odot\mathcal{M}$ ;
6       $\tau\leftarrow\mathcal{P}_{80}(\mathcal{E}[\mathcal{M}>0])$ ;
7       $\mathcal{M}\leftarrow(\mathcal{E}<\tau)\odot\mathcal{M}$ ;
8       $s,b\leftarrow\textbf{Scale Fitting}(\mathbf{D}_{\text{pred}},\mathbf{D}_{\text{sensor}},\mathcal{M})$ ;
9       $\mathbf{D}_{\text{scaled}}\leftarrow s\cdot\mathbf{D}_{\text{pred}}+b$ ;
10
return $\mathbf{D}_{\text{scaled}}$;
Algorithm 1 Dynamic Mask Refinement
Input:
Predicted Depths $\mathbf{D}_{\text{pred}}\in\mathbb{R}^{B\times H\times W}$,
Sensor Depths $\mathbf{D}_{\text{sensor}}\in\mathbb{R}^{B\times H\times W}$,
Mask $\mathcal{M}\in\{0,1\}^{B\times H\times W}$
Output: Scale $s\in\mathbb{R}$, Shift $b\in\mathbb{R}$
1 Flatten and Mask:
$\mathbf{p}$ $\leftarrow\text{Flatten}(\mathbf{D}_{\text{pred}}\odot\mathcal{M})\in\mathbb{R}^{N}$
$\mathbf{s}$ $\leftarrow\text{Flatten}(\mathbf{D}_{\text{sensor}}\odot\mathcal{M})$
@@ 1 Flatten and Mask:
2Construct Global System:
$\begin{bmatrix}\mathbf{p}^{\top}\mathbf{p}&\mathbf{p}^{\top}\mathbf{1}\\ \mathbf{p}^{\top}\mathbf{1}&\mathbf{1}^{\top}\mathbf{1}\end{bmatrix}\begin{bmatrix}s\\ b\end{bmatrix}=\begin{bmatrix}\mathbf{p}^{\top}\mathbf{s}\\ \mathbf{1}^{\top}\mathbf{s}\end{bmatrix}$
Closed-Form Solution:
$s=\frac{(\mathbf{1}^{\top}\mathbf{1})(\mathbf{p}^{\top}\mathbf{s})-(\mathbf{p}^{\top}\mathbf{1})(\mathbf{1}^{\top}\mathbf{s})}{\det}$
$b=\frac{(\mathbf{p}^{\top}\mathbf{p})(\mathbf{1}^{\top}\mathbf{s})-(\mathbf{p}^{\top}\mathbf{1})(\mathbf{p}^{\top}\mathbf{s})}{\det}$
where $\det=(\mathbf{p}^{\top}\mathbf{p})(\mathbf{1}^{\top}\mathbf{1})-(\mathbf{p}^{\top}\mathbf{1})^{2}>\epsilon$
Algorithm 2 Scale Fitting
Figure 10: Depth Scale Alignment: Implementation Details
Normal conditions: For high-precision surface reconstruction, surface normals are critical and can provide surface details rather than only depth. At the same time, normal does not have scale consistency issues. We compute per-frame estimates using the state-of-the-art monocular normal estimator LOTUS [24].
A.3 Appearance conditions
Object Mask: For the manipulation task data, we first select keyframes that effectively capture both the objects and the background conditions of the scene. We then extract all operable objects using a structured prompt template (see Figure 11), which describes each object in terms of its color, material, shape, and spatial position. These object descriptions are fed into Grounding DINO[38] to obtain bounding boxes for object detection. The resulting bounding boxes are subsequently used as prompts for SAM2[49] to generate the corresponding object masks.
*System Instruction*
As a visual descriptor, articulate objects with this flow:
Color -> Material -> Shape/Texture -> Spatial Position
*Template Format*
"A [color] [material] [object], [shape] with [texture], located [region](x,y)"
*Demonstration Examples*
- A red plastic ball, spherical with a smooth glossy surface, at the center (250,300)
- A brown wooden chair, angular with a matte wood grain, in top-left (100,50)
- A silver-blue metallic arm, mechanical with a brushed texture, on the right (600,200)

### v2

@@ 4 Experiments for Synthesis Quality
[…]
Figure 7: Visualizations of RoboTransfer Sim-to-Real generation results.
To evaluate the synthesis quality of RoboTransfer, including multi-view geometric consistency and controllability, we perform both quantitative and qualitative analyses. In this section, we first present the evaluation metrics and analysis of the synthesis quantity in Sec. 4.1. Then, the synthesis quality results are presented in Sec. 4.
4.1 Synthesis Quantitative Analysis
Implementation Details. RoboTransfer is fine-tuned from the pre-trained SVD  [5] model. During inference, we use the EDM scheduler [24] to perform 30 denoising steps and apply classifier-free guidance.
Geometry Consistency Analysis. We first validate the geometric consistency in RoboTransfer’s generative capability. As shown in Table 2, directly using raw depth sensor data (D.S.) as a conditional input introduces considerable noise, which adversely affects the generation quality. This degradation is evident across all three views in terms of depth accuracy, surface normals, and multi-view consistency. In contrast, employing model-predicted [9] depth (D.P.) results in smoother depth maps, which serve as a more stable supervisory signal for training the generative model. Consequently, the generated videos exhibit significantly better geometric consistency. Specifically, compared to D.S.-based generation, D.P.-based conditions relatively improve the RMSE and Mean Err. of the middle view by 27.4% and 14.2%, and increase Pix.Mat. of the middle and left/right views by 4.7% and 37.6%, respectively. Further improvements are observed when we incorporate metric predicted depths (Metric D.P.) to enforce a consistent depth scale across multiple views. This multi-view scale alignment further enhances geometric consistency, leading to relative improvements of 22.4% and 16.8% in Pix.Mat. of left and right views. The Pix.Mat. is omitted for the head camera since it remains static. Finally, combining normal and depth as joint conditions (Metric D.P.+N.) provides complementary geometric cues, yielding the best overall performance across all geometric metrics.
Appearance Consistency Analysis. We then evaluate RoboTransfer’s capability in controlling visual appearance. As shown in Table 3 (row-1 and row-2 of all three camera views), we observe that directly conditioning on background images without inpainting can corrupt the geometric cues. In contrast, inpainting the background before conditioning not only preserves the geometric structure but also improves appearance consistency, leading to $\sim 1\%$ relative improvements in RMSE, median error, and background similarity. Comparing row-1 and row-3 of the head camera view, we examine object-level appearance control. In row-1, the entire image with all the objects is encoded using a global CLIP feature. In contrast, row-3 applies a finer-grained approach by masking and encoding each object individually, followed by feature concatenation. These object-wise conditions lead to a 1% improvement in object CLIP similarity, indicating better control over individual object appearances. Finally, row-4 of all three camera views demonstrates that combining background inpainting with object-wise CLIP encoding achieves the best performance in both foreground and background appearance control. This setup yields the highest CLIP similarity scores for both regions, confirming the effectiveness of jointly modeling structured and unstructured visual components.
4.2 Synthesis Qualitative Result
Real-to-Real Synthesis. We demonstrate that RoboTransfer enables precise, controllable manipulation of scene elements while preserving global geometric and multi-view consistency. As illustrated in Figure 4, our method allows for flexible editing of background attributes (e.g., texture, color) while keeping the foreground and scene geometry intact. Similarly, Figure 5 show that foreground appearance can be altered independently of the background. Figure 6 further highlights the model’s ability to jointly edit both components. By generating diverse variations from structured inputs, this real-to-real framework effectively enriches training data, leading to improved policy generalization for downstream tasks.
Sim-to-Real Synthesis In simulation, per-view geometry conditions can be obtained at no additional cost. As shown in Figure 7, RoboTransfer generates photorealistic videos from simulated geometry inputs, including out-of-distribution cases. This Sim-to-Real paradigm reduces reliance on real-world geometry data, thereby better supporting downstream robotic learning tasks.
Diverse Scene Synthesis For fixed-arm tabletop tasks, backgrounds are simple, while mobile manipulation involves complex geometries. With more training data from [8], Figure 8 shows that RoboTransfer can generate richer, more diverse scenes in complex settings, enhancing the variety and realism of synthetic data for robotic learning.
4.3 Qualitative Comparison
Figure 9 compares robot data generation methods. Cosmos-Transfer [2] performs well under fixed camera views but degrades notably for dynamic viewpoints. In contrast, RoboTransfer maintains strong multi-view consistency, producing realistic and coherent novel-view synthesis. RoboEngine [55], based on image inpainting, suffers from noise and jitter, lacks temporal consistency, and cannot precisely control scene backgrounds or objects.
[…]
Figure 8: Visualizations of RoboTransfer for diversity scene generation results.
[…]
Figure 9: Comparison with Cosmos-Transfer and RoboEngine generation results.
4.4 Comparisons with External Methods
To the best of our knowledge, RoboTransfer is the first method to enable multi-view consistent data synthesis for robotic manipulation. As illustrated in Figure 9, prior approaches suffer from significant multi-view inconsistencies and offer limited background control. Furthermore, standard domain randomization techniques fail to model localized or structurally meaningful variations. In contrast, RoboTransfer leverages a multi-view diffusion model to generate geometry-consistent, high-fidelity variations in both object appearance and background, thereby supporting robust policy learning.
5 Real Robot Experiments
@@
To evaluate the effectiveness of RoboTransfer, we conducted experiments on a real-world robot to assess how synthetic data impacts visual policy generalization.
Table 4: Effectiveness of Synthetic Data Augmentation. Performance comparison across different data augmentation strategies. Success Rate (%) and Score are reported.
Data Composition Spoon Pick&Place Towel Folding
Diff-Obj Diff-All Diff-Obj Diff-All
SR Score SR Score SR Score SR Score
Real only 33.3% 2.67 13.3% 1.56 16.7% 1.81 12% 1.08
Domain Random Aug 44.4% 2.89 11.1% 1.58 16.7% 1.92 12% 1.12
Real + Obj Aug 44.4% 3.00 22.2% 2.04 16.7% 2.00 12% 1.24
Real + Obj&Bg Aug 66.7% 3.56 46.7% 2.98 50.0% 2.60 28% 1.92
5.1 Experimental Platform and Implementation
@@
Experimental Platform and Setup Details. Our experiments were conducted on the Agilex Cobot Magic platform, which is equipped with two PIPER robotic arms and three Intel RealSense D435i cameras, two mounted on the wrists and one positioned overhead. Only RGB data was used for policy learning, presenting a challenging visual understanding problem.
Implementation Details. We adopted ACT [60] as our baseline architecture without modifications. For each task, we collected 100 expert demonstrations via ALOHA [61] teleoperation. To demonstrate generalizability, the scenes and tasks differ from those in the RoboTransfer training data (see Sec. 4 ). More details are provided in the appendix.
5.2 Effectiveness of Synthetic Data
Benchmark Task and Evaluation Methodology. We evaluate synthetic data on two long-horizon dual-arm manipulation tasks: spoon pick-and-place and towel folding. Spoon pick-and-place consists of four phases: the left arm grasps a spoon, places it on a tray, the right arm grasps it, and places it in a cup. Towel folding includes three stages: grasping the bottom corners, lifting and folding upward, and folding the right side to the left. The towel task provides a rigorous test of physical plausibility and temporal coherence in complex manipulations. We introduce a Stage Score to measure phase completion, offering finer-grained insight than binary success metrics and enabling evaluation of policy progression and generalization. Experiments were conducted under two conditions: novel objects (Diff-Obj), and variations in both objects and environment (Diff-All).
Effect of Synthetic Data Proportions. To determine the optimal data mixture, we evaluated policy performance on the spoon pick-and-place task under the Diff-All setting, varying the synthetic data proportion from 0% to 100%. As shown in Figure 10, both metrics peak at a 50/50 ratio. With this balanced distribution, the success rate increases significantly from 13.3% to 46.7%, and the stage score nearly doubles from 1.6 to 3.0. Further increasing the synthetic proportion leads to performance degradation; although geometrically consistent, synthetic data may lack fine-grained physical fidelity (e.g., contact dynamics or material properties), which becomes detrimental at excessive ratios. Notably, a policy trained exclusively on synthetic data still achieves a 40.0% success rate, surpassing the real-only baseline. This highlights the complementary roles of the two sources: synthetic data provides visual diversity, while real data grounds the policy in real-world physical dynamics. Consequently, we adopt the 50/50 ratio for the analysis in the Effectiveness of Synthetic Data section.
@@ 5.2 Effectiveness of Synthetic Data
Figure 10: Performance across synthetic data mixing ratios. Note 0% represents the baseline model trained only on real data.
Effectiveness of Synthetic Data Analysis. As shown in Table 4, our synthetic data augmentation significantly enhances policy robustness where the baseline fails. For the Spoon Pick and Place task under the most difficult Diff-All condition, our approach achieves a 251% relative improvement in success rate (from 13.3% to 46.7%). The results also reveal a clear cumulative benefit of our augmentation strategy. In the Towel Folding Diff-Obj setting, for instance, the Stage Score increases from 1.81 (real only) to 2.0 with object augmentation, and further to 2.6 after adding background augmentation. This confirms that augmenting both object and background appearance is key to learning policies that can handle significant visual domain shifts. While domain randomization handles object color changes, it cannot model localized or structurally meaningful variations.
@@ 6 Conclusion
In this work, we proposed RoboTransfer, a diffusion-based data synthesis framework for robotics that integrates multi-view geometry while providing explicit control over background and object attributes. We also introduced a dataset construction pipeline that generates high-quality triplets incorporating both global geometry and appearance conditions. Experimental results demonstrate that RoboTransfer produces multi-view consistent data, significantly improving the generalization of visuomotor policies for robotic manipulation.
@@
Supplementary Material
7 RoboTransfer Implementation Details
7.1 Training Details
Training Dataset. To construct our training dataset, we leverage the open-source Cobot Magic platform to collect a large-scale video corpus of dual-arm robot executions(see Sec. 8.1 for details). The raw videos are segmented into 10Hz clips of 30 frames each, resulting in approximately 24k clips for training. Additionally, we curate a set of 1.6k 10Hz 30-frame clips from the collected dataset for video synthesis quality evaluation. All videos are annotated with conditions to facilitate both training and assessment.
Training Details. RoboTransfer is fine-tuned from the pre-trained Stable Video Diffusion [5] model. During training, videos from each camera view are resized to a resolution of $640\times 384$. We adopt the AdamW[35] optimizer with a learning rate of $3\times 10^{-5}$ and a global batch size of 8, training for a total of 70K steps. During inference, we use the EDM scheduler [24] to perform 30 denoising steps and apply classifier-free guidance.
7.2 Modeling and Conditions Injection Details
Object Condition. We resize each object image to $224\times 224$ and pass it through CLIP’s image encoder to obtain a single global feature vector per object. These embeddings are then concatenated and fed into the diffusion model. This design allows for finer-grained control, enabling individual manipulation of each object.
Multi-View Consistency Modeling. Previous methods typically introduce a cross-view module to enhance consistency between views. In contrast, RoboTransfer simply concatenates multi-view images, integrating inter-view consistency into global spatial consistency, thereby improving multi-view video modeling. Moreover, this design allows direct loading of pre-trained single-view video generation model weights, without requiring significant modifications to the model architecture.
8 RoboTransfer Dataset Construction Details
For training RoboTransfer, we collected a dedicated dataset and designed a construction pipeline (Figure 3 in the main paper). Here, we provide additional implementation details.
8.1 Data Collection
Robot Platform. We built a large-scale robotic demonstration dataset using the Agilex Cobot Magic platform, following standardized protocols [37, 33]. Each demonstration includes synchronized RGB-D streams from three Intel RealSense D435i cameras: two hand-eye views and one overhead view (Figure 12).
Dataset Design. To capture diverse manipulation scenarios, we designed twelve distinct tasks (Figure 12) with variations in objects, backgrounds, and interactions. For each task, 100 demonstration segments were collected across 10 unique object configurations, totaling 1,000 samples per task. Backgrounds range from textured tabletops to cluttered surfaces, and objects vary in shape, size, and material. This ensures high diversity and realism for robust visual policy training. To further enhance background diversity, we incorporate the AgiBot-World dataset [8].
@@ 8.1 Data Collection
Figure 11: Robot platform visualizations.
@@
Figure 12: Dataset collected with the Agilex robot.
8.2 Geometry Conditions Construction
Depth Conditions. For RGB-D data, raw sensor depth provides accurate metric information but is often noisy or incomplete, particularly on low-reflectivity surfaces. Conversely, simulator-rendered depth is perfect but lacks domain realism. To bridge this gap, we adopt Video Depth Anything (VAD) [9] to generate temporally coherent and spatially complete depth maps. However, VAD predicts relative depth, lacking the global metric scale essential for robotic manipulation. We therefore treat the raw sensor depth as a sparse metric anchor and the VAD output as a dense structural prior. We align them using a robust multi-frame least-squares fitting strategy (Figure 13). Crucially, this process employs an iterative outlier filtering mechanism (Algorithm 1) that allows the model to learn the correct scale $s$ and shift $b$ from reliable sensor pixels while ignoring sensor noise/holes. This ensures the final output inherits the metric accuracy of the sensor while retaining the completeness of VAD. For datasets without multi-view RGB-D sensors, such as AgiBot-World [8], we estimate metric depth using MoGe [48].
Algorithm 1 Dynamic Mask Alignment
@@ 8.2 Geometry Conditions Construction
[…]
3:Initialize mask:
4:   $\mathcal{M}\leftarrow(\mathbf{D}_{\text{sensor}}>\epsilon)\land(\mathbf{D}_{\text{pred}}>\epsilon)$
5:for $i=1$ to 22 do
6:  Scale Fitting:
7:  $s,b\leftarrow\text{ScaleFitting}(\mathbf{D}_{\text{pred}},\mathbf{D}_{\text{sensor}},\mathcal{M})$
8:  $\mathbf{D}_{\text{metric}}\leftarrow s\cdot\mathbf{D}_{\text{pred}}+b$
9:  Mask Update:
10:  $\mathcal{E}\leftarrow|\mathbf{D}_{\text{pred}}-\mathbf{D}_{\text{sensor}}|\odot\mathcal{M}$
11:  $\tau\leftarrow\text{Percentile}_{80}(\mathcal{E}[\mathcal{M}>0])$
12:  $\mathcal{M}\leftarrow(\mathcal{E}<\tau)\odot\mathcal{M}$
13:end for
14:return $\mathbf{D}_{\text{metric}}$
Algorithm 2 Scale Fitting (Least Squares)
[…]
3:Extract valid pixels:
4:   $\mathbf{p}=\mathbf{D}_{\text{pred}}[\mathcal{M}],\quad\mathbf{s}=\mathbf{D}_{\text{sensor}}[\mathcal{M}]$
5:Objective:
6:   $\min_{s,b}\|s\mathbf{p}+b\mathbf{1}-\mathbf{s}\|^{2}$
7:Closed-form Solution:
8:    $\begin{bmatrix}s\\ b\end{bmatrix}=\frac{1}{\Delta}\begin{bmatrix}N(\mathbf{p}^{\top}\mathbf{s})-(\mathbf{1}^{\top}\mathbf{p})(\mathbf{1}^{\top}\mathbf{s})\\ (\mathbf{p}^{\top}\mathbf{p})(\mathbf{1}^{\top}\mathbf{s})-(\mathbf{p}^{\top}\mathbf{s})(\mathbf{1}^{\top}\mathbf{p})\end{bmatrix}$
9:   where $\Delta=N(\mathbf{p}^{\top}\mathbf{p})-(\mathbf{1}^{\top}\mathbf{p})^{2}$
10:return $s,b$
Figure 13: Depth Scale Alignment Algorithms. We employ an iterative robust least-squares fitting strategy. Algorithm 1 aligns relative depth predictions to sparse metric sensor data by iteratively filtering high-error regions (outliers). Algorithm 2 provides the analytical closed-form solution for the linear alignment parameters $s$ (scale) and $b$ (shift).
Normal Conditions Surface normals capture fine geometric details and are scale-invariant. We compute per-frame normals using LOTUS [20]. For datasets lacking multi-view RGB-D sensors, MoGe [48] simultaneously estimates depth and normals, streamlining prelabeling.
8.3 Appearance Conditions Construction
Keyframe Selection. Keyframes serve as the reference for object and background appearance. For standard tabletop tasks, we adopt a temporal strategy: the initial frame (fully populated) defines the object condition, while the final frame (post-manipulation) serves as the background. For complex scenarios, we employ an automated content-aware pipeline using VLM descriptors and Grounding DINO. This pipeline selects the frame maximizing the object’s visible pixel area for the object condition, and conversely, the frame with minimal object presence for the background condition.
Object Mask Generation. Object descriptions, generated via structured prompts (Figure 14) specifying color, material, shape, and spatial position, are fed into Grounding DINO [34] to generate bounding boxes, which SAM2 [41] converts into object masks.
System Prompt for Object Descriptor: You are an industrial robotic vision system with 100% detection guarantee. Perform comprehensive scene analysis to identify ALL movable objects, including occluded items. Strictly enforce size constraints and occlusion handling. Your goal is to complete a list of all visible objects on the table and process the description. The output Template Format: A [color] [object], [shape], located [region](x,y) Examples: A red ball, spherical, at the center (250,300); A brown chair, angular, in the top-left (100,50); A silver-blue arm, mechanical, on the right (600,200);